# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ Abstract
Elliptically symmetric distributions are a classic example of a semiparametric model where the location vector and the scatter matrix (or a parameterization of them) are the two finite-dimensional parameters of interest, while the density generator represents an infinite-dimensional nuisance term. This basic representation of the elliptic model can be made more accurate, rich, and flexible by considering additional finite-dimensional nuisance parameters. Our aim is therefore to investigate the deep and counter-intuitive links between statistical efficiency in estimating the parameters of interest in the presence of both finite and infinite-dimensional nuisance parameters. Unlike previous works that addressed this problem using Le Cam’s asymptotic theory, our approach here is purely geometric: efficiency will be analyzed using tools such as projections and tangent spaces embedded in the relevant Hilbert space. This allows us to obtain original results also for the case where the location vector and the scatter matrix are parameterized by a finite-dimensional vector that can be partitioned in two sub-vectors: one containing the parameters of interest and the other containing the nuisance parameters. As an example, we illustrate how the obtained results can be applied to the well-known “low-rank” parameterization. Furthermore, while the theoretical analysis will be developed for Real Elliptically Symmetric (RES) distributions, we show how to extend our results to the case of Circular and Non-Circular Complex Elliptically Symmetric (C-CES and NC-CES) distributions.
@@ I Introduction
Asemiparametric model is a statistical model that involves not only a finite-dimensional parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ but also an infinite-dimensional parameter, i.e. a function $g$, that often represents a nuisance parameter. This characterization is general enough to include many well-known examples: the symmetric location model, linear and logistic regression, errors in variables models, missing data and censoring models, copula models and even times series models such as ARMA or ARCH models. We refer the reader to [1, Sect. 4] and [2] for a discussion of how the semiparametric formalism applies to the above-mentioned examples. The clear advantage of adopting a semiparametric model is in the potential gain in term of robustness with respect to some missing knowledge about the random experiment at hand that are indeed required when we use a parametric model. On the other hand, the fact that the function $g$ is left fully unspecified may lead to some efficiency losses in the estimation of the parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ in respect of the parametric case. Nevertheless, there are cases in which parametric and semi-parametric efficiency coincide. In other words, in these specific cases, we can gain robustness without losing efficiency. The simplest example is the symmetric location model, in which the goal is to estimate the location parameter without assuming any specific functional form for the (symmetric) density of the data [1, Sect. 3.4, Ex. 1]. In this work, we will analyze the parametric and the semiparametric efficiency for the estimation of the parameter vector of interest when the considered semiparametric model contains nuisance parameters of both finite and infinite dimension. Specifically, we will focus on the statistical model of the elliptically symmetric distributions.
@@
Elliptically symmetric distributions have established themselves as a statistical model capable of capturing the heterogeneous nature of data in a wide range of applications: from remote sensing and communication to seismology and finance (see e.g. [3, 4, 5, 6, 7] for a complete list of references and examples). As for a Gaussian distribution, an elliptical distribution has the advantage of being fully characterized by the location vector $\bm{\mu}$ and the covariance/scatter matrix $\bm{\Sigma}$, while its flexibility with respect to (w.r.t.) the latter is provided by the density generator, i.e. a function $g$, that is able to characterizes the heavy or light-tailed behavior of the data. To this end, the density generator $g$ may depend on additional parameters that control the “level of the tails” of the resulting distribution. Among the most popular and widely-used elliptical distributions, we may cite the Student $t$-distribution (characterized by the, so called, degree of freedom $\lambda$) and the Generalized Gaussian distribution (characterized by the scale and shape parameters) [8].
@@
Inference in elliptical distributions generally requires the estimation of $\bm{\mu}$ and $\bm{\Sigma}$ as main parameters of interest in the eventual presence of additional nuisance terms. For clarity, we would like to recall here that a nuisance term is an unknown parameter whose estimation is not strictly required but that can have an impact on the estimation performance, i.e. the efficiency, of the parameters of interest.
@@
    Parametric modeling with fully specified density generator $g$. This is the less general case in which a full knowledge (i.e. both its functional form and its parameters) of the specific density generator $g$ is assumed to be a-priori known. To fix the idea, in order to derive estimation procedure for $\bm{\mu}$ and $\bm{\Sigma}$, a practitioner may assume to know a-priori that the data follow a $t$-distribution with an a-priori known degree of freedom $\lambda$. This full knowledge of the density generator is not realistic in practice and need to be relaxed.
@@
    Parametric modeling with specified density generator $g$ up to its parameters. This is the classical approach in parametric elliptical inference and most of the literature deals with this case. Returning once again to the example of the $t$-distribution, in this case a practitioner may suppose a-priori that the data generating process is a $t$-distribution characterized by the unknown degree of freedom $\lambda$. Consequently, unlike in M1, the parameters of interest $\bm{\mu}$ and $\bm{\Sigma}$ need to be estimated together with the nuisance term $\lambda$. In general this lack of knowledge on the true value of $\lambda$ may lead to a performance degradation, i.e. efficiency losses, in the estimation of $\bm{\mu}$ and $\bm{\Sigma}$. Even if more flexible w.r.t. the previous modeling approach, the requirement of the a-priori knowledge of the functional form of the density generator $g$ may be questionable. This leads to the semiparametric modeling.
@@
    Semiparametric modeling where the functional form of the density generator $g$ is left unspecified. This is the most realistic case in which the practitioner only supposes that the data are elliptically distributed, without assuming any specific density generator $g$, while estimating the parameters of interest $\bm{\mu}$ and $\bm{\Sigma}$. As a consequence, the density generator has to be considered as an infinite-dimensional nuisance term. Intuitively, one might expect efficiency losses in this case to be greater than those incurred in M2. After all, in model M3, it is the entire functional form of $g$ that is unknown, not just the value of theirs characterizing parameters.
@@
As showed by Hallin and Paindaveine in two seminal papers [9, 10], the answer may be surprising and counter-intuitive. They proved that a decisive role is played by the additional finite-dimensional nuisance parameters that are involved, in an implicit or explicit manner, in the semiparametric modeling of elliptical distributions. Specifically, in their works, Hallin and Paidaveine used the Le Cam’s asymptotic theory to bring to light the fundamental role of a finite-dimensional nuisance parameter hidden in the semiparametric elliptic model: the scale parameter. In short, it is well-known that, in an elliptical model, the scatter matrix $\bm{\Sigma}$ and the density generator $g$ are not jointly identifiable due to a scale ambiguity. In order to remove this ambiguity, $\bm{\Sigma}$ must be rewritten as the product of a scale parameter $s\triangleq S(\bm{\Sigma})$ and of a shape matrix $\mathbf{V}_{S}\triangleq\bm{\Sigma}/S(\bm{\Sigma})$, where $S(\cdot)$ is a given homogeneous functional. In the resulting inference problem, $\bm{\mu}$ and $\mathbf{V}_{S}$ are to be considered as parameters of interest, while $s$ and $g$ are finite and infinite-dimensional nuisance parameters, respectively. Then, in [9, 10], it has been shown that, if the scale parameter $s$ is considered as a nuisance term, then: i) not knowing the function form of $g$ does not lead to any efficiency loss on the estimation of $\bm{\mu}$ and $\mathbf{V}_{S}$ w.r.t. the case in which $g$ is supposed to be a-priori known, and ii) if a determinant-based scale function $S(\cdot)$ is adopted, then knowing $g$ and $s$ does not lead to efficiency losses on $\bm{\mu}$ and $\mathbf{V}_{S}$ w.r.t. the case in which both $g$ and $s$ are perfectly known.
Inspired by this surprising result, in the first part of this paper, we focus on the three models M1, M2 and M3 discussed above (that are slightly different from the scenario analyzed in [9, 10]). Unlike [9, 10] where the main analytical tool is Le Cam’s asymptotic theory, we will use a different approach to the analysis of semiparametric models. Specifically, we will follow the geometric, Hilbert-space-based approach developed in the foundational monograph [1], as it has the potential to reach a broader audience. The main rationale underlying this geometrical approach is that the structure of an Hilbert space enables the unified treatment of both finite-dimensional and infinite-dimensional parameters.
@@
To make this work as self-contained as possible, in Sect. II we present the geometrical tools (i.e. projections and tagent spaces) that will be at the core of our analysis of the semiparametric elliptical model. Moreover, two fundamental lemmas are provided and their proofs detailed in the Appendix. We move then to Sect. III where, after a brief recall of the main definitions and properties of the Real Elliptically Symmetric (RES) distributions, we provide an extensive discussion of score vectors and related Fisher Information Matrix (FIM) for the parametric RES model. In doing this, we will make use of some fundamental outcomes obtained in [9, 10] for the inference of the location $\bm{\mu}$ and the shape $\mathbf{V}_{S}$ in the presence of the nuisance scale parameter $s$. In Sect. IV the geometrical tools introduced in Sect. II will be specified for the semiparametric RES elliptical models under considerations, while our main results of the first part of the paper is collected in Sect. V and in particular in Proposition V.2, where the closed-form expressions of parametric and semiparametric FIM and the related information bounds for the models in M1, M2 and M3 are provided. Proposition V.2 plays a crucial role in formalizing and fully understanding the relationship between the (lack of) efficiency losses and the finite and infinite-dimensional nuisance terms involved in the considered elliptical parametric and semiparametric models.
@@
The second and last part of our paper, basically concentrated in Sect. VI, deals with parameterized elliptical distributions. In particular, we suppose that the location vector $\bm{\mu}(\bm{\theta})$ and the covariance/scatter matrix $\bm{\Sigma}(\bm{\theta})$ are parameterized by a finite dimensional vector $\bm{\theta}=(\bm{\gamma}^{T},\bm{\xi}^{T})^{T}$, where the sub-vector $\bm{\gamma}$ contains the parameters of interest, while $\bm{\xi}$ collects the finite-dimensional nuisance terms. We aim then to investigate the efficiency losses on the following two scenarios:
@@
So we ask ourselves the same question as before: is it possible to relax the parametric assumption, moving from E1 to E2, without losing efficiency? The answer to this question depends on the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ at hand. Proposition VI.1, which represents the main result of Sect. VI, provides a condition to determine whether the given parameterization leads to efficiency losses or not. As an example, this condition will be applied to two cases, important in many applications, that are i) the parameterization $(\bm{\gamma}^{T},\bm{\xi}^{T})^{T}\mapsto(\bm{\mu}(\bm{\gamma}),\bm{\Sigma}(\bm{\xi}))$, i.e. when the location and the scatter matrix have no parameters in common and the ii) “low-rank” parameterization. Then, Sect. VII generalizes all the results previously obtained in the context of RES distributions to the case of Circular Complex Elliptically Symmetric (C-CES) and Non Circular CES (NC-CES) distributions. Finally Sect. VIII concludes the paper and the technical proofs are reported in the Appendix.
@@
$\mathcal{H}\triangleq\left\{h:\mathcal{X}\rightarrow\mathbb{R}|{E_{0}\{h\}}=0,{E_{0}\{h^{2}\}}<+\infty\right\},$ (1)
@@
Remark: It is worth noticing that the $q$-replicating space $\mathcal{H}_{q}$ has been introduced only to simplify some notation and to improve the readability of some result proposed in the following sections. However, it does not introduce any additional geometrical structure since it is just composed of $q$ copies of (,,)($\mathcal{H}$$\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$). In particular, when we work on (subspaces of) $\mathcal{H}$, instead of $\mathcal{H}_{q}$, all the operators, as expectations and projections, has to be interpreted component-wise (see e.g. [13, Sect. 2.4, Remark 2] or [1, Sect. 2.4]) as it is proved in Lemma .1 and Theorem .1 reported in Appendix .1. Specifically, for each $\mathbf{h}=(h_{1},\ldots,h_{q})^{T}\in\mathcal{H}_{q}$ s.t. $h_{i}\in\mathcal{H}$ and $u\in\mathcal{U}\subseteq\mathcal{H}$, we have:
@@ II Models, score vectors and tangent spaces
By using a self-explanatory notation, we now introduce three parametric submodels of $\mathcal{P}$ as:
@@
for $j=1,\ldots,r$. Under the regularity conditions discussed in [17, Sects. 6.2, 6.3], it is immediate to verify that $\mathbf{s}_{\bm{\gamma}_{0}}]_{i},[\mathbf{s}_{\bm{\xi}_{0}}]_{j}\in\mathcal{H}$, $\forall i,j$, i.e. they have zero-mean and finite variance.
We can now introduce the (finite-dimensional) tangent space of the parametric submodels $\mathcal{P}_{2}$ as the linear span of $\mathbf{s}_{\bm{\xi}_{0}}$ in $\mathcal{H}$ [13, Sect. 2.3], [16, eq. (5.24)]:
@@
Finally, the nuisance tangent space $\mathcal{T}_{3}\subseteq\mathcal{H}$ of the non-parametric model $\mathcal{P}_{3}$ is defined as in [1, Sect. 3.2, Def. 2], [13, Sect. 4.4]. Note that, by construction, $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$ are finite- and infinite-dimensional closed subspaces of $\mathcal{H}$.
According to the previous definition, let us define the efficient score vector $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ for the vector $\bm{\gamma}_{0}$ of the parameter of interest in the parametric submodel $\mathcal{P}_{1,2}$ in (8) as [13, Sect. 3.4], [1, Sect. 2.4] and [16, Def. 4]:
@@
where the projection $\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})$ is to be interpreted component-wise as indicated in (3). Since $\mathcal{T}_{2}$ is a finite-dimensional subspace of $\mathcal{H}$, the projection operator $\Pi(\cdot|\mathcal{T}_{2})$ can be derived in closed form as [16, eq. (7)] [13, Sect. 2.4, Ex. 1]:
@@
is the Fisher Information Matrix (FIM) for $\bm{\xi}_{0}$ in the parametric submodel $\mathcal{P}_{2}$ in (7). Then, we have that $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ can be explicitly expressed as:
@@
is the matrix of the cross-information terms in the parametric submodel $\mathcal{P}_{1,2}$ in (8).
@@
for $\bm{\gamma}_{0}$ in the presence of only the finite-dimensional nuisance $\bm{\xi}_{0}$, while $g_{0}$ is known. Specifically, for this comparison, we will make use of the following lemma, given without proof in [1, Sect. 3.4, Prop. 3] and of which we provide a full proof in the Appendix .2 of this work.
@@
It is worth noticing that the matrix $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ is the efficient FIM for $\bm{\gamma}_{0}$ in the presence of the finite-dimensional nuisance vector $\bm{\xi}_{0}$ in the parametric submodel $\mathcal{P}_{1,2}$ in (8). Consequently the Cramér-Rao bound (CRB) for $\bm{\gamma}_{0}$ in the presence of $\bm{\xi}_{0}$ is given by:
@@
is the FIM for $\bm{\gamma}_{0}$ in the parametric submodel $\mathcal{P}_{1}$ in (6). It is immediate to recognize in (26) the well-known expression of the CRB for parametric estimation in the presence of a finite-dimensional nuisance vector.
Moreover, we note that this lemma implies that $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})\leq\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$. This means that the information about the parameter $\bm{\gamma}_{0}$ is reduced or remains the same in the presence of the infinite-dimensional nuisance function $g_{0}$. It is clear from Lemma II.1 that there is no loss of information iff $\mathbf{p}=\mathbf{0}$. The following lemma (also given without proof, in [1, Sect. 3.4, Prop. 3]) and proven in the Appendix .3 of this work, will specify the condition for which the presence of the nuisance function $g_{0}$ will not bring any loss of information.
@@
In the following sections, we will make extensive use of Lemmas II.1 and II.2 to bring to light the sometimes surprising and counter-intuitive relationships between parametric and semiparametric efficiency in the family of elliptical distributions.

### v3

@@ Abstract
Elliptically symmetric distributions are a classic example of a semiparametric model where the location vector and the scatter matrix (or a parameterization of them) are the two finite-dimensional parameters of interest, while the density generator represents an infinite-dimensional nuisance term. This basic representation of the elliptic model can be made more accurate, rich, and flexible by considering additional finite-dimensional nuisance parameters. Our aim is therefore to investigate the deep and counter-intuitive links between statistical efficiency in estimating the parameters of interest in the presence of both finite and infinite-dimensional nuisance parameters. Previous seminal works have addressed this problem by leveraging a general result: if the statistical model has a specific group invariance, then the projection operator onto the semiparametric nuisance tangent space can be asymptotically expressed as a conditional expectation with respect to the maximal invariant sub-$\sigma$ algebra. In this article, we show that, for the statistical model of elliptical distributions, the projection operator can be explicitly computed without relying on the above-mentioned asymptotic approximation. This allows us to obtain original results also for the case in which the location vector and the scatter matrix are parameterized by a finite-dimensional vector that can be partitioned in two sub-vectors: one containing the parameters of interest and the other containing the nuisance parameters. As an example, we illustrate how the obtained results can be applied to the well-known “low-rank” parameterization. Furthermore, while the theoretical analysis will be developed for Real Elliptically Symmetric (RES) distributions, we show how to extend our results to the case of Circular and Non-Circular Complex Elliptically Symmetric (C-CES and NC-CES) distributions.
@@ I Introduction
Asemiparametric model is a statistical model that involves not only a finite-dimensional parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ but also an infinite-dimensional parameter, i.e. a function $g$, that often represents a nuisance parameter. This characterization is general enough to include many well-known examples: the symmetric location model, linear and logistic regression, errors in variables models, missing data and censoring models, copula models and even times series models such as ARMA or ARCH models. We refer the reader to [1, Sect. 4] and [2] for a discussion of how the semiparametric formalism applies to the above-mentioned examples. The clear advantage of adopting a semiparametric model is in the potential gain in term of robustness with respect to some missing knowledge about the random experiment at hand that are indeed required when we use a parametric model. On the other hand, the fact that the function $g$ is left fully unspecified may lead to some efficiency losses in the estimation of the parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ in respect of the parametric case. Nevertheless, there are cases in which parametric and semi-parametric efficiency coincide. In other words, in these specific cases, we can gain robustness without losing efficiency. This property of the statistical model is known as adaptivity [3, 1, 2].
In this work, we will analyze the parametric and the semiparametric efficiency for the estimation of the parameter vector of interest when the considered semiparametric model contains nuisance parameters of both finite and infinite dimension. Specifically, we will focus on the statistical model of the elliptically symmetric distributions.
@@
Elliptically symmetric distributions have established themselves as a statistical model capable of capturing the heterogeneous nature of data in a wide range of applications: from remote sensing and communication to seismology and finance (see e.g. [4, 5, 6, 7, 8] for a complete list of references and examples). As for a Gaussian distribution, an elliptical distribution has the advantage of being fully characterized by the location vector $\bm{\mu}$ and the covariance/scatter matrix $\bm{\Sigma}$, while its flexibility with respect to (w.r.t.) the latter is provided by the density generator, i.e. a function $g$, that is able to characterizes the heavy or light-tailed behavior of the data. To this end, the density generator $g$ may depend on additional parameters that control the “level of the tails” of the resulting distribution. Among the most popular and widely-used elliptical distributions, we may cite the Student $t$-distribution (characterized by the, so called, degree of freedom $\nu$) and the Generalized Gaussian distribution (characterized by the scale and shape parameters) [9].
@@
Inference in elliptical distributions generally requires the estimation of $\bm{\mu}$ and $\bm{\Sigma}$ as main parameters of interest in the eventual presence of additional nuisance terms. For clarity, we recall that a nuisance term is an unknown parameter whose estimation is not strictly required but which can affect the performance of the estimator of the parameter of interest, i.e. the efficiency, of the parameters of interest.
@@
    Parametric modeling with fully specified density generator $g$. This is the least general case in which a full knowledge (i.e. both its functional form and its parameters) of the specific density generator $g$ is assumed to be a-priori known. To fix the idea, in order to derive estimation procedure for $\bm{\mu}$ and $\bm{\Sigma}$, a practitioner may assume to know a-priori that the data follow a $t$-distribution with an a-priori known degree of freedom $\nu$. This full knowledge of the density generator is not realistic in practice and need to be relaxed.
@@
    Parametric modeling with specified density generator $g$ up to its parameters. This is the classical approach in parametric elliptical inference and most of the literature deals with this case. Returning once again to the example of the $t$-distribution, in this case a practitioner may suppose a-priori that the data generating process is a $t$-distribution characterized by the unknown degree of freedom $\nu$. Consequently, unlike in M1, the parameters of interest $\bm{\mu}$ and $\bm{\Sigma}$ need to be estimated together with the nuisance term $\nu$. In general this lack of knowledge on the true value of $\nu$ may lead to a performance degradation, i.e. efficiency losses, in the estimation of $\bm{\mu}$ and $\bm{\Sigma}$. Even if more flexible w.r.t. the previous modeling approach, the requirement of the a-priori knowledge of the functional form of the density generator $g$ may be questionable. This leads to the semiparametric modeling.
@@
    Semiparametric modeling where the functional form of the density generator $g$ is left unspecified. This is the most realistic case in which the practitioner only supposes that the data are elliptically distributed, without assuming any specific density generator $g$, while estimating the parameters of interest $\bm{\mu}$ and $\bm{\Sigma}$. As a consequence, the density generator has to be considered as an infinite-dimensional nuisance term. Intuitively, one might expect efficiency losses in this case to be greater than those incurred in M2. After all, in model M3, it is the entire functional form of $g$ that is unknown, not just the value of its characterizing parameters.
@@
As showed by Hallin and Paindaveine in two seminal papers [10, 11], the answer may be surprising and counter-intuitive. They proved that a decisive role is played by the additional finite-dimensional nuisance parameters that are involved, in an implicit or explicit manner, in the semiparametric modeling of elliptical distributions. Specifically, in their works, Hallin and Paidaveine built upon a general result, proposed in [2, Prop. 2.6], to bring to light the fundamental role of a finite-dimensional nuisance parameter hidden in the semiparametric elliptic model: the scale parameter. In short, it is well-known that, in an elliptical model, the scatter matrix $\bm{\Sigma}$ and the density generator $g$ are not jointly identifiable due to a scale ambiguity. In order to remove this ambiguity, $\bm{\Sigma}$ must be rewritten as the product of a scale parameter $s\triangleq S(\bm{\Sigma})$ and of a shape matrix $\mathbf{V}_{S}\triangleq\bm{\Sigma}/S(\bm{\Sigma})$, where $S(\cdot)$ is a given homogeneous functional of order one. In the resulting inference problem, $\bm{\mu}$ and $\mathbf{V}_{S}$ are to be considered as parameters of interest, while $s$ and $g$ are finite and infinite-dimensional nuisance parameters, respectively. Then, in [10, 11], it has been shown that, if the scale parameter $s$ is considered as a nuisance term, then: i) not knowing the functional form of $g$ does not lead to any efficiency loss on the estimation of $\bm{\mu}$ and $\mathbf{V}_{S}$ w.r.t. the case in which $g$ is a-priori known (so-called restricted adaptivity [10]), and ii) if a determinant-based scale function $S(\cdot)$ is adopted, then not knowing $g$ and $s$ does not lead to efficiency losses on $\bm{\mu}$ and $\mathbf{V}_{S}$ w.r.t. the case in which both $g$ and $s$ are perfectly known (the full adaptivity property).
Inspired by this surprising result, in the first part of this paper, we focus on the three models M1, M2 and M3 discussed above. Specifically, our goal here is to recover the adaptivity conditions found in [10, 11], but using a more direct approach. In fact, the approach proposed in [10, 11] is to derive this condition of adaptivity from a more general result proposed in [2, Prop. 2.6]. Without going into detail, this general result shows that, for certain models, the projection operator can be asymptotically approximated as a conditional expectation with respect to the maximal invariant (with respect to the group that generates the statistical model) sub-$\sigma$ algebra. In particular, the generating group of the RES distribution, along with its maximal invariant statistics, has been introduced and discussed in [12]. In contrast, in this article we show how to derive, for the elliptical model, this projection operator can be derived directly without resorting to its asymptotic approximation.
@@
To make this work as self-contained as possible, in Sect. II we present the geometrical tools (i.e. projections and tangent spaces) that will be at the core of our analysis of the semiparametric elliptical model. Moreover, two fundamental lemmas are provided and their proofs detailed in the Appendix. We move then to Sect. III where, after briefly recalling of the main definitions and properties of the Real Elliptically Symmetric (RES) distributions, we provide an extensive discussion of score vectors and related Fisher Information Matrix (FIM) for the parametric RES model. In doing this, we will make use of some fundamental outcomes obtained in [10, 11] for the inference of the location $\bm{\mu}$ and the shape $\mathbf{V}_{S}$ in the presence of the nuisance scale parameter $s$. In Sect. IV the geometrical tools introduced in Sect. II will be specified for the semiparametric RES elliptical models under considerations, while the main results of the first part of the paper are collected in Sect. V, in particular in Proposition V.2, where the closed-form expressions of parametric and semiparametric FIM and the related information bounds for the models in M1, M2 and M3. Proposition V.2 plays a crucial role in formalizing and fully understanding the relationship between the (lack of) efficiency losses and the finite and infinite-dimensional nuisance terms involved in the considered elliptical parametric and semiparametric models.
@@
The second and last part of our paper, basically concentrated in Sect. VI, deals with parametrized elliptical distributions. In particular, we suppose that the location vector $\bm{\mu}(\bm{\theta})$ and the covariance/scatter matrix $\bm{\Sigma}(\bm{\theta})$ are parameterized by a finite dimensional vector $\bm{\theta}=(\bm{\gamma}^{T},\bm{\xi}^{T})^{T}$, where the sub-vector $\bm{\gamma}$ contains the parameters of interest, while $\bm{\xi}$ collects the finite-dimensional nuisance terms. We aim then to investigate the efficiency losses on the following two scenarios:
@@
So we ask ourselves the same question as before: is it possible to relax the parametric assumption, moving from E1 to E2, without losing efficiency? The answer to this question depends on the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ at hand. Proposition VI.1, which represents the main result of Sect. VI, provides an adaptivity condition that allows one to determine whether the given parameterization leads to efficiency losses or not. As an example, this condition will be applied to two cases of practical importance: i) the parameterization $(\bm{\gamma}^{T},\bm{\xi}^{T})^{T}\mapsto(\bm{\mu}(\bm{\gamma}),\bm{\Sigma}(\bm{\xi}))$, i.e. when the location and the scatter matrix have no parameters in common and the ii) “low-rank” parameterization. Then, Sect. VII generalizes all the results previously obtained in the context of RES distributions to the case of Circular Complex Elliptically Symmetric (C-CES) and Non Circular CES (NC-CES) distributions. Finally Sect. VIII concludes the paper and the technical proofs are reported in the Appendix.
@@
$\mathcal{H}\triangleq\left\{h\in L_{2}(P_{0})|{E_{0}\{h\}}=0\right\}=\left\{h:\mathcal{X}\rightarrow\mathbb{R}|{E_{0}\{h\}}=0,{E_{0}\{h^{2}\}}<+\infty\right\},$ (1)
@@
Remark: It is worth noticing that the $q$-replicating space $\mathcal{H}_{q}$ has been introduced only to simplify some notation and to improve the readability of some result proposed in the following sections. However, it does not introduce any additional geometrical structure since it is just composed of $q$ copies of (,,)($\mathcal{H}$$\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$). In particular, when we work on (subspaces of) $\mathcal{H}$, instead of $\mathcal{H}_{q}$, all the operators, such as expectations and projections, has to be interpreted component-wise (see e.g. [15, Sect. 2.4, Remark 2] or [1, Sect. 2.4]) as is proved in Lemma .1 and Theorem .1 reported in Appendix .1. Specifically, for each $\mathbf{h}=(h_{1},\ldots,h_{q})^{T}\in\mathcal{H}_{q}$ s.t. $h_{i}\in\mathcal{H}$ and $u\in\mathcal{U}\subseteq\mathcal{H}$, we have:
@@ II Models, score vectors and tangent spaces
By using a self-explanatory notation, we now introduce three parametric sub-models of $\mathcal{P}$ as:
@@
for $j=1,\ldots,r$. Under the regularity conditions discussed in [19, Sects. 6.2, 6.3], it is immediate to verify that $\mathbf{s}_{\bm{\gamma}_{0}}]_{i},[\mathbf{s}_{\bm{\xi}_{0}}]_{j}\in\mathcal{H}$, $\forall i,j$, i.e. they have zero-mean and finite variance.
We can now introduce the (finite-dimensional) tangent space of the parametric sub-models $\mathcal{P}_{2}$ as the linear span of $\mathbf{s}_{\bm{\xi}_{0}}$ in $\mathcal{H}$ [15, Sect. 2.3], [18, eq. (5.24)]:
@@
Finally, the nuisance tangent space $\mathcal{T}_{3}\subseteq\mathcal{H}$ of the non-parametric model $\mathcal{P}_{3}$ is defined as in [1, Sect. 3.2, Def. 2], [15, Sect. 4.4]. Note that, by construction, $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$ are finite- and infinite-dimensional closed subspaces of $\mathcal{H}$.
According to the previous definition, let us define the efficient score vector $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ for the vector $\bm{\gamma}_{0}$ of the parameter of interest in the parametric sub-model $\mathcal{P}_{1,2}$ in (8) as [15, Sect. 3.4], [1, Sect. 2.4] and [18, Def. 4]:
@@
where the projection $\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})$ is to be interpreted component-wise as indicated in (3). Since $\mathcal{T}_{2}$ is a finite-dimensional subspace of $\mathcal{H}$, the projection operator $\Pi(\cdot|\mathcal{T}_{2})$ can be derived in closed form as [18, eq. (5.7)] [15, Sect. 2.4, Ex. 1]:
@@
is the Fisher Information Matrix (FIM) for $\bm{\xi}_{0}$ in the parametric sub-model $\mathcal{P}_{2}$ in (7). Then, we have that $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ can be explicitly expressed as:
@@
is the matrix of the cross-information terms in the parametric sub-model $\mathcal{P}_{1,2}$ in (8).
@@
for $\bm{\gamma}_{0}$ in the presence of only the finite-dimensional nuisance $\bm{\xi}_{0}$, while $g_{0}$ is known. Specifically, for this comparison, we will make use of the following lemma, stated without proof in [1, Sect. 3.4, Prop. 3], for which we provide a full proof in the Appendix .2 of this paper.
@@
It is worth noticing that the matrix $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ is the efficient FIM for $\bm{\gamma}_{0}$ in the presence of the finite-dimensional nuisance vector $\bm{\xi}_{0}$ in the parametric sub-model $\mathcal{P}_{1,2}$ in (8). Consequently the Cramér-Rao bound (CRB) for $\bm{\gamma}_{0}$ in the presence of $\bm{\xi}_{0}$ is given by:
@@
is the FIM for $\bm{\gamma}_{0}$ in the parametric sub-model $\mathcal{P}_{1}$ in (6). It is immediate to recognize in (26) the well-known expression of the CRB for parametric estimation in the presence of a finite-dimensional nuisance vector.
Moreover, we note that this lemma implies that $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})\leq\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$. This means that the information about the parameter $\bm{\gamma}_{0}$ is reduced or remains the same in the presence of the infinite-dimensional nuisance function $g_{0}$. It is clear from Lemma II.1 that there is no loss of information iff $\mathbf{p}=\mathbf{0}$. The following lemma (also stated without proof, in [1, Sect. 3.4, Prop. 3]) and proven in the Appendix .3 of this work, specifies the condition under which the presence of the nuisance function $g_{0}$ will not bring any loss of information.
@@
In Sect. IV and in particular in Sect. VI, we will make extensive use of this framework and of Lemma II.2 to bring to light the sometimes surprising and counter-intuitive relationships between parametric and semiparametric efficiency in the family of elliptical distributions.